Abstracts of talks

Speaker: Francisco Garcia Date: 07/09/2009 Title: Strict convexity on non-complete spaces

Abstract: We will construct a non-strictly convex Banach space with a dense strictly convex subspace.

Speaker: Mrinal Raghupathi Date: 07/13/2009 Title: A Nevanlinna-Pick theorem for multiplier algebras

Abstract: Given a set X, a kernel K on X, points x_{1}, …, x_{n} ∈ X, and scalars w_{1}, …, w_{n} ∈ ℂ, the Nevanlinna-Pick problem is to determine conditions for the existence of a function in the multiplier algebra of K such that f(x_{j}) = w_{j} for j = 1, …, n.

In this talk we will review some recent results in Nevanlinna-Pick interpolation. We will show how the one-step extension method of Agler, McCullough and Quiggin can be applied to prove a Nevanlinna-Pick type theorem for the multiplier algebra of a kernel function on a set X. This result contains as special cases the known interpolation theorems for subalgebras of H^{∞}.

Speaker: Daniel Freeman Date: 07/14/2009 Title: A greedy basis for (∑ l_{p}^{n})_{l_{q}}

Abstract: A basis (x_{i})_{i=1}^{∞} for a Banach space X is called greedy if for every x = ∑_{i=1}^{∞} a_{i}x_{i} ∈ S_{X} the best n-term approximation of x is up to a constant given by taking the n terms with the largest coeficients in absolute value. Greedy bases were characterized by Konyagin and Temlyakov to be exactly those bases which are both democratic and unconditional. We prove that the Banach space (∑_{n=1}^{∞} ℓ_{p}^{n})_{ℓ_{q}} has a greedy basis whenever 1 ≤ p ≤ ∞ and 1 < q < ∞. Furthermore, we prove that the Banach space (∑_{n=1}^{∞} ℓ_{p}^{n})_{ℓ_{1}} never has a greedy basis whenever 1 < p ≤ ∞. This is joint work with Stephen Dilworth, Edward Odell, and Thomas Schlumprecht.

Speaker: Yun-Su Kim Date: 07/17/2009 Title: Hilbert Spaces with respect to Operator-Valued Norms

Abstract: In this presentation, we introduce two kinds of operator- valued norms. One of them is an L(H)-valued norm. The other one is an L(C(K))-valued norm. We provide the notion of a Hilbert space with respect to an L(H)-valued norm. By a Hilbert space with respect to an L(H)-valued norm, we generalize the notion of a Hilbert space. Furthermore, we provide several interesting examples (L^{∞} and C(Y)) of Hilbert spaces with respect to an operator-valued norm.

Speaker: Florent Baudier Date: 07/27/2009 Title: A new metric invariant

Abstract: In 1986, J. Bourgain characterized the superreflexivity of a Banach space X in terms of bi-Lipschitz embedding of hyperbolic dyadic trees (B_{n}, ρ). This result is a concrete occurrence of the ”the Ribe program”. Bourgain’s theorem can be restated as follows : D(X) > ω (or D(X^{∗}) > ω) if and only if every (B_{n}, ρ) bi-Lipschitz embeds into X with a universal positive constant distortion C, where ω is the first infinite countable ordinal and D(X) denotes the dentability index of X. This is also equivalent to the embedding of the infinite dyadic tree. The talk will focus on a similar result when dealing with the relation between Szlenk index and countably branching trees. Applications concerning stability results for certain classes of Banach spaces will be given. This is a joint work with Nigel Kalton and Gilles Lancien.

Speaker: Ping Wong Ng Date: 07/28/2009 Title: Projection Decomposition in Operator Algebra

Abstract: Motivated by the work of Dykema, Freeman, Kornelson, Larson, Ordower and Weber in frame theory, we study when a positive operator in a C^{∗}-algebra can be written as a (possibly infinite) sum of (not necessarily pairwise orthogonal) projections.

For type I and type III von Neumann factors (with separable predual), we have complete characterizations. (Here, the sums of projections converge in the strong operator topology.) If B is a sigma-unital simple stable purely infinite C^{∗}-algebra then for the multiplier algebra M(B) of B, we have a complete characterization. (Here, the sums of projections converge in the strict topology.) For other cases, we have partial results. This is joint work with V. Kaftal and S. Zhang.

Speaker: Bentuo Zheng Date: 07/29/2009 Title: On the unconditional subsequence property

Abstract: We show that a construction of Johnson, Maurey and Schechtman leads to the existence of a weakly null sequence (f_{i}) in (∑ L_{p_{i}})_{ℓ_{2}}, where p_{i} ↓ 1, so that for all ε > 0 and 1 < q ≤ 2, every subsequence of (f_{i}) admits a block basis (1+ε)-equivalent to the Haar basis for L_{q}. We give an example of a reflexive Banach space having the unconditional subsequence property but not uniformly so.

Speaker: Nirina Randrianarivony Date: 07/30/2009 Title: Asymptotic uniform smoothness and nonlinear embeddings

Abstract: Abstract: This talk complements Florent Baudier’s talk last Monday by presenting the chronological state of knowledge before the Baudier, Kalton and Lancien’s result. In this talk, we will see how the asymptotic uniform smoothness of a Banach space X is used to study the structure of another space that can coarse bilipschitz embed into it. This is joint work with Nigel Kalton.

Speaker: Kevin Beanland Date: 08/10/2009 Title: Weak Hilbert Spaces with Few Symmetries

Abstract: In this talk we will give an outline of the construction of a weak Hilbert space 𝔛_{wh} with an unconditional basis and having the property that every operator on a block subspace Y of 𝔛_{wh} can be written as a diagonal operator plus a strictly singular operator. This implies that no block subspace of 𝔛_{wh} is isomorphic to any of its proper subspaces. We will also present possible modifications of this construction and discuss the difficulties involved in constructing a weak Hilbert space exhibiting more conditional structure. The work we present here is joint with S.A. Argyros and Th. Raikoftsalis.

Speaker: Timur Oikhberg Date: 08/11/2009 Title: Bernstein’s “lethargy” in the non-commutative setting

Abstract: A classical theorem of S.Bernstein states that, for any increasing sequence of finite dimensional subspaces E_{1} ↪ E_{2} ↪ … ↪ X (X is a Banach space), and for any sequence α_{i} ↘ 0, there exists x ∈ X s.t. dist(x, E_{i}) = α_{i} for every i. In this talk, we consider certain non-commutative analogues of this results. We prove that, for any pair of infinite dimensional Banach spaces X and Y, and for any sequence α_{i} ↘ 0, there exists T ∈ B(X, Y) whose sequences of approximation, Gelfand, and Kolmogorov numbers “behave like” (α_{i}). Other s-scales are also considered.

Speaker: Francesco Fidaleo Date: 08/12/2009 Title: Ergodic properties of Bogoliubov automorphisms in free probability

Abstract: We show that some C^{∗}–dynamical systems obtained by ”quantizing” classical ones on the free Fock space, enjoy very strong ergodic properties. Namely, if the classical dynamical system (X, T, μ) is ergodic but not weakly mixing, then the resulting quantized system (𝔊, α) is uniquely ergodic (w.r.t the fixed point algebra) but not uniquely weak mixing. The same happens if we quantize a classical system (X, T, μ) which is weakly mixing but not mixing. In this case, the quantized system is uniquely weak mixing but not uniquely mixing. Finally, a quantized system arising from a classical mixing dynamical system, will be uniquely mixing. In such a way, it is possible to exhibit uniquely weak mixing and uniquely mixing C^{∗}–dynamical systems whose GNS representation associated to the unique invariant state generates a von Neuman factor of one of the following types: I_{∞}, II_{1}, III_{λ} where λ ∈ (0, 1].